Problem 4.13 Question 11 of 14 Check My Work (No more tries available) Click here to read the eBook: Debt Management Ratios TIE AND ROIC RATIOS The W.C. Pruett Corp. has $800,000 of interest-bearing debt outstanding, and it pays an annual interest rate of 9%. In addition, it has $600,000 of common stock on its balance sheet. It finances with only debt and common equity, so it has no preferred stock. Its annual sales are $2 million, its average tax rate is 35%, and its profit margin is 8%, what are its TIE ratio and its return on invested capital (ROIC)? Round your answers to two decimal places. ROIC Hide Feedback Incorrect Check My Work (No more tries available) Icon Key Problem 4.13 Question 11 of 14
Answer #1

Earnings after tax = Sales * profit margin = $2 million * 8% = $160,000

If EAT is $160,000 and tax rate is 35%, Earnings before tax is =160000/(1-0.35) = 246,153.84

Interest = outstanding amount * int. rate = $800,000 * 9% = $72,000

If EBT is 246,153.84 and interest is 72,000, then EBIT is sum of EBT and int = 246153.84 + 72000 = $318,153.85

a) TIE ratio is EBIT/Int. = 318153.85/72000 = 4.42

b) ROIC = Net income / invested capital(debt+equity)

=246,153.84 / 1,400,000

=17.58%